|5-3x+|3x+1||-4=-2x
Answer by lwsshak3(11628)

|5-3x+|3x+1||-4=-2x
**
check first possible true condition: expressions within the bars are>0, so just remove the bars and solve like a regular equation:
5-3x+3x+1-4=-2x
2x=-2
x=-1
check: |5-3x+|3x+1||-4=-2x
=|5+3+|-3+1||-4+2x=0
=|5+3-3+1|-4-2
=6-4-2=0
..
check second possible true condition: expressions within the bars are<0, so place a negative sign in front of the expressions before removing the bars then solve like a regular equation:
|5-3x+|3x+1||-4=-2x
-[5-3x-(3x+1)]-4=-2x
-[5-3x-3x-1]-4=-2x
-5+3x+3x+1-4=-2x
6x-8=-2x
8x=8
x=1
Check:
|5-3x+|3x+1||-4=-2x
=|5-3+4|-4+2x=0
=6-4+2=4≠0
reject, x=1 is not a solution
x=-1 is the only solution